ID |
Date |
Author |
Type |
Category |
Subject |
15690
|
Wed Nov 25 18:30:23 2020 |
gautam | Update | ASC | Some thoughts about AS WFS electronics | An 8 channel whitening chassis was prepared and tested. I measured:
- TF from input to output - there are 7 switchable stages (3 dB, 6 dB, 12 dB and 24 dB flat whitening gain, and 3 stages of 15:150 Hz z:p whitening). I enabled one at a time and measured the TF.
- Noise with input terminated.
In summary,
- All the TFs look good (I will post the plots later), except that the 3rd stage of whitening on both boards don't show the expected transfer function. The fact that it's there on both boards makes me suspect that the switching isn't happening correctly (I'm using a little breakout board). I'm inclined to not debug this because it's unlikely we will ever use 3 stages of 15:150 whitening for the AS WFS.
- The noise measurement displayed huge (x1000 above the surrounding broadband noise floor) 60 Hz harmonics out to several kHz. My hypothesis is that this has to do with some bad grounding. I found that the circuit ground is shorted to the chassis via the shell of the 9pin and 15pin Dsub connectors (but the two D37 connector shields are isolated). This seems very wierd, idk what to make of this. Is this expected? Looking at the schematic, it would appear that the shields of the connectors are shorted to ground which seems like a bad idea. afaik, we are using the same connectors as on the chassis at the sites - is this a problem there too? Any thoughts?
Quote: |
Whitening chassis. Waiting for front panels to arrive, PCBs and interface board are in hand, stuffed and ready to go. A question here is how we want to control the whitening - it's going to be rather difficult to have fast switchable whitening. I think we can just fix the whitening state. Another option would be to control the whitening using Acromag BIO channels.
|
|
15689
|
Wed Nov 25 18:18:41 2020 |
gautam | Update | ASC | Planned mods for WFS head | I am confused by the discussion during the call today. I revisited Hartmut's paper - the circuit in Fig 6 is essentially what I am calling "only 2f_2 notch stuffed" in my previous elog. Qualitatively, the plot I presented in Attachment #2 of the preceeding elog in this thread shows the expected behavior as in Fig 8 of the paper - the impedance seen by the photodiode is indeed lower. In Attachment #1, I show the comparison - the "V(anode)/I(I1)" curve is analogous to the "PD anode" curve in Hartmut's paper, and the "V(vout)/I(I1)" curve is analogous to the "1f-out" curve. I also plot the sensitivity analysis (Attachment #2), by varying the photodiode junction capacitance between 100pF and 200 pF (both values inclusive) in 20 pF steps. There is some variation at 55 MHz, but it is unlikely that the capacitance will change so much during normal operation?
I understand the motivation behind stuffing the other notches, to reduce intermodulation effects. But the impression I got from the call was that somehow, the model I presented was wrong. Can someone help me identify the mistake?
I didn't bother to export the LTspice data and make a matplotlib plot for this quick analysis, so pardon the poor presentation. The colors run from green=100pF to grey=200pF. |
Attachment 1: anodeVsOutput.png
|
|
Attachment 2: sensitivity.png
|
|
15688
|
Tue Nov 24 16:51:29 2020 |
gautam | Update | PonderSqueeze | Ponderomotive squeezing in aLIGO | Summary:
On the call last week, I claimed that there isn't much hope of directly measuring Ponderomotive Squeezing in aLIGO without some significant configurational changes. Here, I attempt to quantify this statement a bit, and explicitly state what I mean by "significant configurational changes".
Optomechanical coupling:
The I/O relations will generally look something like:
.
The. magnitudes of the matrix elements C_12 and C_21 (i.e. phase to amplitude and amplitude to phase coupling coefficients) will encode the strength of the Ponderomotive squeezing.
Readout:
For the inital study, let's assume DC readout (since there isn't a homodyne readout yet even in Advanced LIGO). This amounts to setting in the I/O relations, where the former angle is the "homodyne phase" and the latter is the "SRC detuning". For DC readout, the LO quadrature is fixed relative to the signal - for example, in the usual RSE operation, . So the quadrature we will read out will be purely (or nearly so, for small detunings around RSE operation). The displacement noises will couple in via the matrix element. Attachment #1 and Attachment #2 show the off-diagonal elements of the "C" matrix for detunings of the SRC near RSE and SR operation respectively. You can see that the optomechanical coupling decays pretty rapidly above ~40 Hz.
SRC detuning:
In this particular case, there is no benefit to detuning the SRC, because we are assuming the homodyne angle is fixed, which is not an unreasonable assumption as the quadrature of the LO light is fixed relative to the signal in DC readout (not sure what the residual fluctuation in this quantity is). But presumably it is at the mrad level, so the pollution due to the orthogonal anti-squeezed quadrture can be ignored for a first pass I think. I also assume ~10 degrees of detuning is possible with the Finesse ~15 SRC, as the linewidth is ~12 degrees.
Noise budget:
To see how this would look in an actual measurement, I took the data from Lee's ponderomotive squeezing paper, as an estimate for the classical noises, and plotted the quantum noise models for a few representative SRC detunings near RSE operation - see Attachment #3. The curves labelled for various phis are the quantum noise models for those SRC detunings, assuming DC readout. I fudged the power into the IFO to make my modelled quantum noise curve at RSE line up with the high frequency part of the "Measured DARM" curve. To measure Ponderomotive Squeezing unambiguously, we need the quantum noise curve to "dip" as is seen around 40 Hz for an SRC tuning of 80 degrees, and that to be the dominant noise source. Evidently, this is not the case.
The case for balanced homodyne readout:
I haven't analyzed it in detail yet - but it may be possible that if we can access other quadratures, we might benefit from rotating away from the DARM quadrature - the strength of the optomechanical coupling would decrease, as demonstrated in Attachments #1 and #2, but the coupling of classical noise would be reduced as well, so we may be able to win overall. I'll briefly investigate whether a robust measurement can be made at the site once the BHD is implemented. |
Attachment 1: QN_heatmap_RSE.pdf
|
|
Attachment 2: QN_heatmap_SR.pdf
|
|
Attachment 3: noiseBudget.pdf
|
|
15687
|
Mon Nov 23 23:27:43 2020 |
Koji | Summary | ASC | Q3000 characterization | Last week and this week I've been working on the characterization of the Q3000 QPDs. The QPDs were named 81, 82, 83, and 94.
- Dark current [OMC LAB ELOG 402]: All the segments looked similar and acceptable except for the seg1 of #82. It has a smaller reverse breakdown voltage (~6V) but even this is an acceptable level.
- Impedance [OMC LAB ELOG 403]: All the segments showed a ~300pF junction capacitance with no reverse bias. This looks quite normal.
- Dark noise [OMC LAB ELOG 404]: All the segments showed ~5pA/rtHz dark noise above 1Hz.
My recommendation is to use #81 and #84 as they have similar dark current characteristics between the segments. But basically, all the QPDs look fine.
The actual junction capacitance and the RF dark noise should be characterized by the actual WFS head circuit.
The QPD packages were labeled and returned to Gautam to be implemented in the WFS heads.
gautam: S/N #84 was installed as the AS WFS QPD. The remaining 3 are stored in the clean cabinet at EX (where the rest of the RF photodiodes are). |
15686
|
Mon Nov 23 16:33:10 2020 |
gautam | Update | VAC | More vacuum deliveries | Five Agilent pressure gauges were delivered to the 40m. It is stored with the controller and cables in the office area. This completes the inventory for the gauge replacement - we have all the ordered parts in hand (though. not necessarily all the adaptor flanges etc). I'll see if I can find some cabinet space in the VEA to store these, the clutter is getting out of hand again...
in addition, the spare gate valve from LHO was also delivered today to the 40m. It is stored at EX with the other spare valves.
Quote: |
It is stored along with the cables that arrived a few weeks ago, awaiting the gauges which are now expected next week sometime.
|
|
15685
|
Mon Nov 23 14:52:10 2020 |
Koji | Update | BHD | BHD MMT Optics delivered | Basically, they repeated our specs and showed the coating performances for HR/AR for 10deg P and PR/AR for 45deg P. There is no RoC measurement by the vendor.
Nevertheless, their RoC (paper) specs should be compared with our request. |
15684
|
Mon Nov 23 12:25:14 2020 |
gautam | Update | BHD | BHD MMT Optics delivered | Optics --> Cabinet at south end (Attachment #1)
Scanned datasheets--> wiki. It would be good if someone can check the specs against what was ordered. |
Attachment 1: IMG_8965.HEIC
|
15683
|
Sun Nov 22 21:09:37 2020 |
gautam | Update | ASC | Planned mods for WFS head | Attachment #1 - Proposed mods for 40m RF freqs.
- I followed Rich's suggestion of choosing an inductor that has Z~100 ohms at the frequency of interest.
- The capacitor is then chosen to have the correct resonant frequency.
- Voltronix trim caps are used for fine tuning the resonances. 2 variants are used, one with a range of 4-20 pF, and a Q of 500 per spec, while the other has range of 8-40 pF, and a Q of 200 per spec.
- In the table, the first capacitance is the fixed one, and the second is the variable one. We're not close to the rail for the variable caps.
- For the first trials, I think we can try by not populating all of the notches - just the 2f notch. We can then add notches if deemed necessary. Probably these notches are more important for a REFL/POP port WFS.
- One thing I noticed is that the aLIGO WFS use ceramic capacitors for the LC reactances. i haven't checked if there is any penalty we are paying in terms of Q of the capacitor. anyways, i'm not going to redesign the PCB and maybe ceramic is the only option in the 0805 package size?
Attachment #2 - Modelled TFs for the case where all the notches are stuffed, and where only the 2f notch is stuffed.
- The model uses realistic composite models for the inductors from coilcraft, but the capacitors are idealized parts.
- I also found the library part for LMH6624, so this should be a bit closer to the actual circuit than Rich's models which subbed in the MAX4107 in place.
- The dashed vertical lines indicate some frequencies of interest.
- Approx 1 kohm transimpedance is realized at 55 MHz. I don't have the W/rad number for the sensitivity at the AS port, but my guess is this will be just fine.
- If the 44 MHz and 66 MHz notches are stuffed, then there is some interaction with the 55 MHz notch, which lowers the transimpedance gain somewhat. So if we decide to stuff those notches, we should do a mroe careful investigation into whether this is problematic.
Attachment #3 - Modelled TFs for the case where all the notches are stuffed, and where only the 2f notch is stuffed.
- Initially, I found the (modelled) noise level to be rather higher than expected. It persisted despite making the resistors in the model noiseless. Turns out there is some leakage from the "Test Input" path. Some documents in the DCC suggest that there should be an "RF Relay" that allows one to isolate this path, but afaik, the aLIGO WFS does not have this feature. So maybe what we should do is to remove C9 once we're done tuning the resonances. Better yet, just tune the resonance with the Jenne laser and not this current-injection path.
- Horizontal dashed lines indicate shot noise for the indicated DC photocurrent levels. It is unlikely we will have even 1 mW of light on a single quadrant at the AS port, so the AS port WFS will not be shot noise limited. But I think that's okay for initial trials.
- The noise level of ~20 pA/rtHz input referred is in agreement what I would expect using Eq 3 of the LMH6624 datasheet. The preamp has a gain of 10, so the source impedance seen by it is ~100 ohms (since the overall gain is 1kohm). The corresponding noise level per Eq 3 is ~2 nV/rtHz, or 20 pA/rtHz current noise referred to the photocurrent 👍 .
- The LMH6624 datasheet claims that the OpAmp is stable for CLG >= 10. For reasons that aren't obvious to me, Koji states here that the CLG needs to be even higher, 15-20 for stability. Do the aLIGO WFS see some instability? Should I raise R14 to 900 ohms?
Any other red flags anyone sees before I finish stuffing the board?
Quote: |
WFS head and housing. Need to finalize the RF transimpedance gain (i.e. the LC resonant part), and also decide which notches we want to stuff.
|
|
Attachment 1: aLIGO_wfs_v5_40m.pdf
|
|
Attachment 2: TFs.pdf
|
|
Attachment 3: noise.pdf
|
|
15682
|
Wed Nov 18 22:49:06 2020 |
gautam | Update | ASC | Some thoughts about AS WFS electronics | Where do we want to install the interface and readout electronics for the AS port WFS? Options are:
- 1Y1 / 1Y3 (i.e. adjacent to the LSC rack) - advantage is that 55 MHz RF signal is readily available for demodulation. But space is limited (1Y2, where the RF signal is, is too full so at the very least, we'd have to run a short cable to an adjacent rack), and we'd have a whole bunch of IPC channels between c1lsc and c1ioo models.
- 1X1/1X2. There's much more space and we can directly digitize into the c1ioo model, but we'd have to route the 55 MHz signal back to this rack (kind of lame since the signal generation is happening here). I'm leaning towards this option though - thinking we can just open up the freq generation box and take a pickoff of the 55 MHz signal...
There isn't much difference in terms of cable length that will be required - I believe the AS WFS is going to go on the AP table even in the new optical layout and not on the ITMY in-air oplev table?
The project requires a large number of new electronics modules. Here is a short update and some questions I had:
- WFS head and housing. Need to finalize the RF transimpedance gain (i.e. the LC resonant part), and also decide which notches we want to stuff. Rich's advise was to not stuff any more than is absolutely necessary, so perhaps we can have at first just the 2f notch and add others as we deem necessary once we look at the spectrum with the interferometer locked. Need to also figure out a neat connector solution to get the signals from the SMP connectors on the circuit board to the housing - I'm thinking of using Front-Panel-Express to design a little patch board that we can use for this purpose, I'll post a more detailed note about the design once I have it.
- WFS interface board + soft-start board (the latter provides a smooth ramp up of the PD bias voltage). These go in a chassis, the assembly is almost complete, just waiting on the soft-start board from JLCPCB. One question is how to power this board - Sorensens or linear? If we choose to install in 1X1/1X2, I guess Sorensen is the only option, unless we have a couple of linear power supplies lying around spare.
- Demod board (quad chassis). Assembly is almost complete, need to install the 4 way RF splitter, some insulating shoulder washers. (to ensure the RF ground is isolated from the chassis), and better nuts for the D-sub connectors. A related question is how we want to supply the electrical LO signal for demodulation. The "nominal" level each demod board wants is 10 dBm. This signal will be sourced inside the chassis from a 4-way RF splitter (~7 dB insertion loss). So we'd need 17dBm going into the splitter. This is a little too high for a compact amplifier like the ZHL-500-HLN to drive (1dB compression point is 16 dBm), and the signal level available at the LSC rack is only ~2 dBm. So do we want a beefy amplifier outside the chassis amplifying the signal to this level? Or do we want to use the ZHL-500-HLN, and amplify the signal to, say 13 dBm, and drive each board with ~6 dBm LO? The Peregrine mixer on these boards (PE4140) are supposed to be pretty forgiving in terms of the LO level they want... In either case, I think we should avoid having an amplifier also inside the chassis, it is rather full in there with 4 demod boards, regulator board, all the cabling, and an RF splitter. It may be that heat dissipation becomes an issue if we stick an RF amplifier in there too...
- Whitening chassis. Waiting for front panels to arrive, PCBs and interface board are in hand, stuffed and ready to go. A question here is how we want to control the whitening - it's going to be rather difficult to have fast switchable whitening. I think we can just fix the whitening state. Another option would be to control the whitening using Acromag BIO channels.
- AI chassis - will go between whitening and ADC.
- Large number of cables to interconnect all the above pieces. I've asked Chub to order the usual "Deluxe" shielded Dsub cables, and we will get some long SMA-SMA cables to transmit the RF signals from head to demod board from Pasternack (or similar), do we need to use Heliax or the Times Microwave alternative for this purpose? What about the LO signal? Do we want to use any special cable to route it from the LSC rack to the IOO rack, if we end up going that way?
Approximately half of the assembly of the various electronics is now complete. The basic electrical testing of the interface chassis and demod chassis are also done (i.e. they get power, the LEDs light up, and are stable for a few minutes). Detailed noise and TF characterization will have to be done. |
15681
|
Wed Nov 18 17:51:50 2020 |
gautam | Update | VAC | Agilent pressure gauge controller delivered | It is stored along with the cables that arrived a few weeks ago, awaiting the gauges which are now expected next week sometime. |
15680
|
Tue Nov 17 13:24:40 2020 |
Chub | Update | General | big UPS on the way | Ordered 11/16 from CDW, on PO# S492940, the high voltage Tripp Lite SMART5000XFMRXL for TP-1. Should be arriving in about a week. |
15679
|
Tue Nov 17 00:26:32 2020 |
rana | Update | Electronics | SR560s in need of repair/battery replacement | yes, both problems can be fixed. Usually we just order some spare lead-acid batteries from SRS (Steve may have some spare ones somewhere). The DC offset often comes from a busted FET input. I bought 50 of those at one point - they're obsolete. Its also possible to replace the input stage with any old FET pair.
I'll handle the one with the offset if you leave it on my desk. |
15678
|
Mon Nov 16 16:00:19 2020 |
gautam | Update | Equipment loan | LB1005-->Cryo lab | Shruti picked it up @4pm. |
15677
|
Mon Nov 16 00:02:34 2020 |
rana | Update | Electronics | More systematic noise characterization | true. also try to choose a cap with a goow high frequency response. In the Electronics Noise book by Ott there's some graph about this. I bet you good do a Bing search and also find something more modern. Basically we want to make sure that the self resonance is not happening at low frequencies. Might be tought to find one with a good HF response, a high voltage rating, and > 1uF.
Quote: |
Yes. The datasheet has a recommendation circuit with 10uF caps. Companies are careful to show reproducible, reliably functional circuit examples on datasheets. So, if the caps are there you should try to replicate the design.
Quote: |
Additional bypass capacitors? I use 0.1 uF, 700V DC ceramic capacitors as bypass capacitors close to the leads of the PA95, as is recommended in the datasheet. Can adding a 10uF capacitor in parallel provide better filtering? I'm not sure if one with compatible footprint and voltage rating is readily available, I'll look around.
|
|
|
15676
|
Thu Nov 12 15:40:42 2020 |
Koji | Update | Electronics | More systematic noise characterization | Yes. The datasheet has a recommendation circuit with 10uF caps. Companies are careful to show reproducible, reliably functional circuit examples on datasheets. So, if the caps are there you should try to replicate the design.
Quote: |
Additional bypass capacitors? I use 0.1 uF, 700V DC ceramic capacitors as bypass capacitors close to the leads of the PA95, as is recommended in the datasheet. Can adding a 10uF capacitor in parallel provide better filtering? I'm not sure if one with compatible footprint and voltage rating is readily available, I'll look around.
|
|
15675
|
Thu Nov 12 14:55:35 2020 |
gautam | Update | Electronics | More systematic noise characterization | Summary:
I now think the excess noise in this circuit could be coming from the KEPCO switching power supply (in fact, the supplies are linear, and specd for a voltage ripple at the level of <0.002% of the output - this is pretty good I think, hard to find much better).
Details:
All component references are w.r.t. the schematic. For this test, I decided to stuff a fresh channel on the board, with new components, just to rule out some funky behavior of the channel I had already stuffed. I decoupled the HV amplifier stage and the Acromag DAC noise filtering stages by leaving R3 open. Then, I shorted the non-inverting input of the PA95 (i.e. TP3) to GND, with a jumper cable. Then I measured the noise at TP5, using the AC coupling pomona box (although in principle, there is no need for this as the DC voltage should be zero, but I opted to use it just in case). The characteristic bump in the spectra at ~100Hz-1kHz was still evident, see the bottom row of Attachment #1. The expected voltage noise in this configuration, according to my SPICE model, is ~10 nV/rtHz, see the analysis note.
As a second test, I decided to measure the voltage noise of the power supply - there isn't a convenient monitor point on the circuit to directly probe the +/- HV supply rails (I didn't want any exposed HV conductors on the PCB) - so I measured the voltage noise at the 3-pin connector supplying power to the 2U chassis (i.e. the circuit itself was disconnected for this measurement, I'm measuring the noise of the supply itself). The output is supposedly differential - so I used the SR785 input "Float" mode, and used the Pomona AC coupling box once again to block the large DC voltage and avoid damage to the SR785. The results are summarized in the top row of Attachment #1.
The shape of the spectra suggests to me that the power supply noise is polluting the output noise - Koji suggested measuring the coherence between the channels, I'll try and do this in a safe way but I'm hesitant to use hacky clips for the High Voltage. The PA95 datasheet says nothing about its PSRR, and seems like the Spice model doesn't include it either. It would seem that a PSRR of <60dB at 100 Hz would explain the excess noise seen in the output. Typically, for other Op-Amps, the PSRR falls off as 1/f. The CMRR (which is distinct from the PSRR) is spec'd at 98 dB at DC, and for other OpAmps, I've seen that the CMRR is typically higher than the PSRR. I'm trying to make a case here that it's not unreasonable if the PA95 has a PSRR <= 60dB @100 Hz.
So what are the possible coupling mechanisms and how can we mitigate it?
- Use better power supply - I'm not sure how this spec of 10-50 uV/rtHz from the power supply lines up in the general scheme of things, is this already very good? Or can a linear power supply deliver better performance? Assuming the PSRR at 100 Hz is 60 dB and falls off as 1/f, we'd need a supply that is ~10x quieter at all frequencies if this is indeed the mechanism.
- Better grounding? To deliver the bipolar voltage rails, I used two unipolar supplies. The outputs are supposedly floating, so I connected the "-" input of the +300 V supply to the "+" input of the -300 V supply. I think this is the right thing to do, but maybe this is somehow polluting the measurement?
- Additional bypass capacitors? I use 0.1 uF, 700V DC ceramic capacitors as bypass capacitors close to the leads of the PA95, as is recommended in the datasheet. Can adding a 10uF capacitor in parallel provide better filtering? I'm not sure if one with compatible footprint and voltage rating is readily available, I'll look around.
What do the analog electronics experts think? I may be completely off the rails and imagining things here.
Update 2130: I measured the coherence between the positive supply rail and the output, under the same conditions (i.e. HV stage isolated, input shorted to ground). See Attachment #2 - the coherence does mirror the "bump" seen in the output voltage noise - but the coherence is. only 0.1, even with 100 averages, suggesting the coupling is not directly linear - anyways, I think it's worth it to try adding some extra decoupling, I'm sourcing the HV 10uF capacitors now. |
Attachment 1: powerSupplyNoise.pdf
|
|
Attachment 2: coherence.pdf
|
|
15674
|
Thu Nov 12 14:31:27 2020 |
gautam | Update | Electronics | SR560s in need of repair/battery replacement | I had to go through five SR560s in the lab yesterday evening to find one that had the expected 4 nV/rtHz input noise and worked on battery power. To confirm that the batteries were charged, I left 4 of them plugged in overnight. Today, I confirmed that the little indicator light on the back is in "Maintain" and not "Charge". However, when I unplug the power cord, they immediately turn off.
One of the units has a large DC output offset voltage even when the input is terminated (though it is not present with the input itself set to "GND" rather than DC/AC). Do we want to send this in for repair? Can we replace the batteries ourselves? |
Attachment 1: IMG_8947.jpg
|
|
15673
|
Thu Nov 12 14:26:35 2020 |
gautam | Update | General | ETMY suspension eigenmodes | The results from the ringdown are attached - in summary:
- The peak positions have shifted <50 mHz from their in-air locations, so that's good I guess
- The fitted Qs of the POS and SIDE eigenmodes are ~500, but those for PIT and YAW are only ~200
- The fitting might be sub-optimal due to spurious sideband lobes around the peaks themselves - I didn't go too deep into investigating this, especially since the damping seems to work okay for now
- There is up to a factor of 5 variation in the response at the eigenfrequencies in the various sensors - this seems rather large
- The condition number of the matrix that would diagonalize the sensing is a scarcely believable 240, but this is unsurprising given the large variation in the response in the different sensors. Unclear what the implications are - I'm not messing with the input matrix for now
|
Attachment 1: ETMY.tar.bz2
|
15672
|
Tue Nov 10 17:46:06 2020 |
gautam | Update | General | IFO recovery | Summary:
- Recovery was complicated by RFM failure on c1iscey - see Attachment #1. This is becoming uncomfortably frequent. As a result, the ETMY suspension wasn't being damped properly. Needed a reboot of c1iscey machine and a restart of the c1rfm model to fix.
- POX/POY locking was restored. Arm alignment was tuned using the dither alignment system.
- AS beam was centered on its CCD (I put a total of ND=1.0 filters back on the CCD). Note that the power in the AS beam is ~4x what it was given we have removed the in-vacuum pickoff to the OMC.
- Green beams were aligned to the arm cavities. See Attachment #2. Both green cameras were adjusted on the PSL table to have the beam be ~centered on them.
- ALS noise is far too high for locking, needs debugging. See Attachment #3.
- AS beam was aligned onto the AS55 photodiode. With the PRM aligned, the REFL beam was centerd on the various REFL photodiodes. The PRMI (resonant carrier) could be locked, see Attachment #4.
I want to test out an AS port WFS now that I have all the parts in hand - I guess the Michelson / PRMI will suffice until I make the ALS noise good again, and anyways, there is much assembly work to be done. Overnight, I'm repeating the suspension eigenmode measurement. |
Attachment 1: RFMerrs.png
|
|
Attachment 2: IFOrecovery.png
|
|
Attachment 3: ALS_ool.pdf
|
|
Attachment 4: PRMIcarr.png
|
|
15671
|
Tue Nov 10 15:13:41 2020 |
rana | Update | General | ETMY suspension eigenmodes | For the input matrix diagonalization, it seemed to me that when we had a significant seismic event or a re-alignment of the optic with the bias sliders, the input matrix also changes.
Meaning that our half-light voltage may not correspond to the half point inside the LED beam, but that rather we may be putting the magnet into a partially occluding state. It would be good to check this out by moving the bias to another setting and doing the ringdown there. |
15670
|
Tue Nov 10 14:30:06 2020 |
gautam | Update | IOO | WFS2 broken | While proceeding with the interferometer recovery, I noticed that there appeared to be no light on WFS2. I confirmed on the AP table that the beam was indeed hitting the QPD, but the DC quadrants are all returning 0. Looking back, it appears that the failure happened on Monday 26 October at ~6pm local time. For now, I hand-aligned the IMC and centered the beams on the WFS1 and MC2T QPDs - MCT is ~15000 cts and MC REFL DC is ~0.1, all consistent with the best numbers I've been able to obtain in the past. I don't think the servo will work without 1 sensor without some retuning of the output matrix.
It would appear that both the DC and RF outputs of WFS2 are affected - I dithered the MC2 optic in pitch (with the WFS loop disabled) at 3.33 Hz, the transmission and WFS1 sensors see the dither but not WFS2. It could be that I'm just not well centerd on the PD, but by eye, I am, so it would appear that the problem is present in both the DC and RF signal paths. I am not going into the PD head debugging today.
Quote: |
Looking back through the elog, 1mtorr is the pressure at which it is deemed safe to send the full power beam into the IMC. After replacing the HR mirror in the MCREFL path with a 10% reflective BS, I just cranked the power back up. IMC is locked. With the increased exposure on the MC2T camera, lots of new scattered light has become visible.
|
|
Attachment 1: WFS2broken.png
|
|
Attachment 2: WFS2broken_RF.png
|
|
15669
|
Tue Nov 10 12:41:23 2020 |
gautam | Update | IOO | 1W —> IMC | Looking back through the elog, 1mtorr is the pressure at which it is deemed safe to send the full power beam into the IMC. After replacing the HR mirror in the MCREFL path with a 10% reflective BS, I just cranked the power back up. IMC is locked. With the increased exposure on the MC2T camera, lots of new scattered light has become visible. |
Attachment 1: CD93A725-FB5C-4F67-BB2E-D122205114B0.jpeg
|
|
15668
|
Tue Nov 10 11:59:37 2020 |
gautam | Update | VAC | Stuck RV2 | I've uploaded some more photos here. I believe the problem is a worn out thread where the main rotary handle attaches to the shaft that operates the valve.
This morning, I changed the valve config such that TP2 backs TP1 and that combo continues to pump on the main volume through the partially open RV2. TP3 was reconfigured to pump the annuli - initially, I backed it with the AUX drypump but since the load has decreased now, I am turning the AUX drypump off. At some point, if we want to try it, we can try pumping the main volume via the RGA line using TP2/TP3 and see if that allows us to get to a lower pressure, but for now, I think this is a suitable configuration to continue the IFO work.
There was a suggestion at the meeting that the saturation of the main volume pressure at 1mtorr could be due to a leak - to test, I closed V1 for ~5 hours and saw the pressure increased by 1.5 mtorr, which is in line with our estimates from the past. So I think we can discount that possibility. |
Attachment 1: damagedThread.001.jpeg
|
|
Attachment 2: IFOstatus.png
|
|
Attachment 3: P1a_leakTest.png
|
|
15667
|
Tue Nov 10 11:31:13 2020 |
Koji | Update | General | Pumpdown | Main volume pressure as of 11:30AM 2020/11/10 |
Attachment 1: Screen_Shot_2020-11-10_at_11.30.21.png
|
|
15666
|
Mon Nov 9 17:40:17 2020 |
gautam | Update | General | Pumpdown | [koji, rana, gautam]
1100 - EY chamber inspected, no issues were found --> EY heavy door on
1200 - OMC chamber was inspected. OM6 was marginally tweaked to bring the beam down a little in pitch, and also a little northwards in Yaw. --> Heavy door on.
1230 - Pumpdown started. Initially, the annuli volume was pumped down. The procedure calls for doing this with the small turbopumps. However, V7 was left open, and hence, in the process, the TP1 foreline pressure (=P2) hit ~30 torr. This caused TP1 to shutdown. We were able to restart it without issue. This case was not caught by the interlock code, which was running at the time. It should be recitified.
1330 - OMC breadboard clean optics and DCPD hardware were wrapped up and packed into tupperware boxes and stored along the south arm. OMC cavity itself, the OMMT, and the breadboard the OMC was sitting on are wrapped in foil/Ameristat and stored in cabinet S13, lower 2 shelves.
1915 - P1a = 0.5 torr pressure reached. Switched over to pumping the main volume with TP1, backed by TP2 and TP3, which themselves are backed by their respective dry pumps and also the AUX drypump for some extra oomph. All cooling fans available in the area were turned on and directed at the turbo pumps. RV2 was used to throttle the flow suitably.
It was at this point that we hit a snag - RV2 has gotten stuck in a partially open position, see Attachment #1. We can see that the thread doesn't move in response to turning the rotary dial. Fortunately, the valve is partially open, so the main volume continues to be pumped - see Attachment #2 for the full history of today's pumping. We are leaving the main volume pumped in this configuration overnight (TP1 pumping main volume backed by TPs 2 and 3, which are in turn backed by their respective drypumps and also the AUX dry pump). I think there is little to no risk of any damage to the turbo pumps, the interlocks should catch any anomalies. The roughing pumps RP1 and RP3 were turned off and that line was disconnected and capped.
What are our options?
- The main volume is able to reach the "nominal" pressure of 1e-5 torr, just takes longer.
- At some point, we may be able to pump the main volume directly with TP2 and TP3 - it is unclear at this point whether it's better to have the conductance limited TP1, which has a higher pumping capacity, or have the smaller TPs 2 and 3 pump through a larger conductance.
- Depending on how low the ultimate pressure gets, we may be able to run the usual IFO activities until the replacement pressure gauges arrive in ~1 week, at which point we can vent the pumpspool (leaving the main volume isolated) and either repair this valve or replace it with one of the spares we have.
We need some vacuum experts to comment. Why did this happen? Is this an acceptable failure mode of the valve?
KA Ed:
2230 - P1a = 0.025 torr. The pressure is coming down with log-linear scale. x0.1 per 2.5 hours or so. |
Attachment 1: IMG_8937.jpg
|
|
Attachment 2: pumpdown.png
|
|
15665
|
Sun Nov 8 23:24:29 2020 |
Koji | Update | General | Vent activities today | > I didn't bother to align the green beams to the arm cavities or re-center the Oplevs - is this necessary? It is a step in the pre-close up checklist, so maybe we should do it... The green transmission does reach the PSL table...
I don't think so. The beam is reaching the PSL, so we have no motivation to change the green alignment. Regarding the oplev, the green refl should come back to the PDH PD and this gives us additional beam reference. As soon as we find the green resonance after the pumping, we can tweak the green axis so that the spots on the mirrors become reasonable (as well as the green trans CCD on the PSL table).
|
15664
|
Sun Nov 8 20:22:35 2020 |
gautam | Update | General | Vent activities today | Basic IFO alignment checks were done.
- IMC could be locked - I tweaked the cavity alignment a little to maximize the MC transmission.
- Y arm can be made resonant for a TEM00 mode of the main beam. I can't run ASS successfully in this low power config. I can see that the axis isn't great, the spot is visually off-center on ITMY, but we should have plenty of actuator range to correct for this with TT1/TT2.
- X arm shows IR mode flashes in the TRX QPD. The green beam can be made resonant for a TEM00 mode, but that alignment doesn't yield the largest IR resonant peaks in TRX. I suspect it is due to the mis-alignment of the beam axis.
- AS beam was aligned onto the CCD. I could see clean Michelson fringes by tweaking the BS alignment. On the AP table, I noticed that the beam on the first steering mirror after the AS beam exits the vacuum is a little high. We can easily resolve this by tweaking OM6 pitch a bit, but even if we don't I don't see any major issues as there is plenty of clearance w.r.t. the viewport when the beam exits the vacuum.
- With the PRM aligned, I can see the REFL beam on the CCD, and it doesn't look clipped.
- I didn't bother to align the green beams to the arm cavities or re-center the Oplevs - is this necessary? It is a step in the pre-close up checklist, so maybe we should do it... The green transmission does reach the PSL table...
Tomorrow, we should do some visual checks of the chambers / EQ stops on ETMY etc but I don't see any major problems at the moment...
Quote: |
Barring any catastrophic failures and provided all required personnel are available, we will do the final pre-close-up checks, put the heavy doors back on, and pump down starting 10 am Monday, 9 Nov 2020.
|
|
Attachment 1: IFOalignment.png
|
|
15663
|
Fri Nov 6 14:27:16 2020 |
gautam | Update | CDS | c1bhd setup - diskless boot | I was able to boot one of the 3 new Supermicro machines, which I christened c1bhd, in a diskless way (with the boot image hosted on fb, as is the case for all the other realtime FEs in the lab). This is just a first test, but it is reassuring that we can get this custom linux kernel to boot on the new hardware. Some errors about dolphin drivers are thrown at startup but this is to be expected since the server isn't connected to the dolphin network yet. We have the Dolphin adaptor card in hand, but since we have to get another PCIe card (supposedly from LLO according to the BHD spreadsheet), I defer installing this in the server chassis until we have all the necessary hardware on hand.
I also have to figure out the correct BIOS settings for this to really run effectively as a FE (we have to disable all the "un-necessary" system level services) - these machines have BIOS v3.2 as opposed to the older vintages for which there are instructions from K.T. et al.
There may yet be issues with drivers, but this is all the testing that can be done without getting an expansion chassis. After the vent and recovering the IFO, I may try experimenting with the c1ioo chassis, but I'd much prefer if we can do the testing offline on a subnet that doesn't mess with the regular IFO operation (until we need to test the IPC).
Quote: |
I am working on the setup of a CDS FE, so please do not attempt any remote login to the IPMI interface of c1bhd until I'm done.
|
|
Attachment 1: Screenshot_2020-11-06_14-26-54.png
|
|
15662
|
Fri Nov 6 14:08:44 2020 |
gautam | Update | PSL | PMC re-locked | The PMC servo railed and so I re-locked it at ~half range. I've been noticing that the diurnal drift of the PZT control voltage has been larger than usual - not sure if it's entirely correlated with temperature on the PSL table. Anyway the cavity is locked again so all is good. |
15661
|
Fri Nov 6 11:36:37 2020 |
gautam | Update | General | ETMY suspension eigenmodes | Attachment #1 shows the main result - there are 4 peaks. The frequencies are a little different from what I have on file for ETMY and the Qs are a factor of 3-4 lower (except SIDE) than what they are in vacuum, which is not unreasonable I hypothesize. The fits suggest that the peak shape isn't really Lorentzian, the true shape seems to have narrower tails than a Lorentzian, but around the actual peak, the fit is pretty good. More detailed diagnostic plots (e.g. coil-to-coil TFs) are in the compressed Attachment #2. The condition number of the matrix to diagonalize the sensing matrix (i.e. what we multiply the "naive" OSEM 2 Euler basis matrix by) is ~40, which is large, but I wouldn't read too much into it at this point.
I see no red flags here - the PIT peak is a little less prominent than the others, but looking back through the elog, this kind of variation in peak heights doesn't seem unreasonable to me. If anyone wants to look at the data, the suspension was kicked every ~1100seconds from 1288673974, 15 times.
Quote: |
I'm measuring the free-swinging spectra of ETMY overnight.
|
|
Attachment 1: ETMY_pkFitNaive.pdf
|
|
Attachment 2: ETMY.tar.bz2
|
15660
|
Thu Nov 5 22:27:34 2020 |
gautam | Update | General | Vent activities today | [koji, gautam]
- Tweaked last steering mirror before PSL beam is launched into vacuum to get the IMC resonant for the TEM00 mode. Less than 1/8 turn in each Pitch and Yaw was required, and we recovered MC2T ~ 1300 cts (c.f. the expected ~1500 cts when the cavity is well aligned, but we didn't touch the IMC mirrors).
- ETMY magnet re-gluing
- Regluing was successful. Pickle picker came off smoothly.
- We performed the razor blade test on all 6 magnets, the integrity of the joints seems uncompromised.
- ETMY cleaning
- Applied F.C. to HR face, HR edge (to remove some residual stains) and AR surface (taking care to go around the magnets).
- A fresh bottle each of spectroscopic grade acetone and isopropanol was dispensed into clean beakers for the work.
- Curiously, in the lighting conditions of the cleanroom, both HR and AR faces looked surprisingly clean, even when illuminated with the green flashlight.
- However, when we took it to the east end and looked at it again with the green flashlight, just before putting it into the chamber, we saw all sorts of stains and markings on the HR side.
- The central aperture looks fine - the contamination did not happen during the transport so we feel confident not to futz around more with this and since we cleaned it extensively in the cleanroom, we opted not to do another round of cleaning in the chamber.
- ETMY re-suspension (on the existing wire loop)
- The back OSEM-holding plates were removed for access reasons.
- The optic was rested onto the EQ stops.
- Patient nudging of the wire around the optic allowed us to eventually get the wire into the V-grooves in the standoffs.
- The wire was not damaged! (at least, to eye, free-swinging test is going on now, we will see what the eigenfrequencies and Qs are)...
- We confirmed that the wire is seated in the v-groove, and took some close up photos.
- ETMY pitch balancing
- A HeNe was brought into the cleanroom and mounted at 5.5" beam height.
- Level-ness of the beam (i.e. parallel to the optic table) was done coarsely with a spirit level placed on the HeNe holder, and more fine adjustment was done by checking the beam height at the launch point, and again ~2 meters away.
- All EQ stops were backed off to ensure the optic was really "free-swinging" - no OSEMs either at this point, since all this was done in the cleanroom.
- The return beam hit the HeNe's output aperture, so we were satisfied that the pitch balancing was good to ~1mrad.
- ETMY transport back to EY
- Since we want to maintain some tension in the wires, the bottom EQ stops were not raised as much as they were when the tower was transported from EY to the cleanroom.
- As a result, we felt it'd be better for a human (Koji) to carry the cage (rather than cart it, the idea being that there would be fewer vertical impulses due to bumps on the ground etc).
- The transport went smooth.
- EY chamber work
- Optic was placed back in its original position (marked by some L-clamps).
- In preparation for alignment work, I undid all the CDS changes I made to facilitate the temporary 3-magnet actuation. This necessitated a reboot of c1auxey VME crate.
- OSEMs were inserted - best effort to half light as usual, and we tried to replicate the rotation in the mounts as closely as possible to what was the case before (using a photo from the last vent as reference).
- Coarse alignment was done by making the Oplev return beam hit the QPD.
- Better alignment of the Y-arm cavity axis was done using the green beam.
- IMC was locked - the beam alignment target was used to ensure the IR beam was hitting ETMY at the correct height (by moving TT2), and then the ETMY angular alignment was improved to make the return beam go back roughly collinearly.
- Y-arm POY locking
- Satisfied with the alignment, we returned to the control room (light doors back on EY chamber) - cavity alignment was tweaked further to make the TRY resonances as high as possible.
- I could lock the Y-arm length to the PSL frequency using POY as a sensor - TRY was ~0.04-0.05 (since IMCT is 1/10th its usual value, we expect it to be 0.1, so this is not bad).
- OMC chamber work
- Tied down the balance weights that were previously placed after the OMC and peripheral optics were removed.
- Checked the table leveling.
- Checked that the AS beam is reasonably well centerd on OM5 and OM6. Took some photos.
- Checked that the IMC could be locked after this work - it could.
So all the primary vent objectives have been achieved 🙌 . The light doors are on the chamber right now. I'm measuring the free-swinging spectra of ETMY overnight. Barring any catastrophic failures and provided all required personnel are available, we will do the final pre-close-up checks, put the heavy doors back on, and pump down starting 10 am Monday, 9 Nov 2020. Some photos here. |
15659
|
Wed Nov 4 17:14:49 2020 |
gautam | Update | CDS | c1bhd setup | I am working on the setup of a CDS FE, so please do not attempt any remote login to the IPMI interface of c1bhd until I'm done.
|
15658
|
Wed Nov 4 00:10:43 2020 |
gautam | Update | General | Vent activities today | [koji, gautam]
- Glued broken off magnet - curing overnight with lamp to slightly elevate temp for curing.
- Remvoed material from OMC chamber as per the plan. This is all sitting wrapped up in foil on a cart for tonight, we should figure out a better storage plan eventually.
The IMC isn't resonant for a TEM00 mode at the time of writing - we are waiting for the stack to relax, at which point if the IMC isn't resonant for a TEM00 mode, we will tweak the input pointing into the IMC (we want to use the suspended cavity as the reference, since it is presumably more reliable than the table from which we removed ~50 kgs of weight and shifted the balance. |
Attachment 1: IMG_0184.jpg
|
|
Attachment 2: IMG_0185.jpg
|
|
Attachment 3: IMG_0187.jpg
|
|
Attachment 4: IMG_0207.jpeg
|
|
Attachment 5: IMG_0210.jpg
|
|
15657
|
Tue Nov 3 09:06:50 2020 |
gautam | Update | General | Lab alarm tripped | I got a call from Calum ~830am today saying some facilities people entered the lab, opened the south entrance door, and tripped the alarm in the process. I came to the lab shortly after and was able to reset the alarm by flipping the switch on the alarm box at the south end entrance to "Alarm OFF". Then, I double checked that the door is closed, and re-enabled the alarm. The particle count at the SP table is not unusually high and the lasers (Oplev HeNe and AUX X) were still on, so doesn't look like any lasting damage was done. The facilities people were apparently wearing laser safety goggles. |
15656
|
Mon Nov 2 17:32:05 2020 |
gautam | Update | General | Plan for OMC chamber | Good point - looking back, I also see that I already removed the mirror at the SW corner of the table in 2016. Revised photo in Attachment #1. There is an optic on the east edge of this table whose purpose I'm not sure of, but I'm pretty sure it isn't essential to the main functionality and so can be removed.
Quote: |
I believe the mirror next to IM1 is for the green beams to be delivered to the PSL table. I think we still want to keep it. Otherwise, the plan looks fine.
|
|
Attachment 1: IMG_2317.JPG
|
|
15655
|
Mon Nov 2 17:13:19 2020 |
Koji | Update | General | Plan for OMC chamber | I believe the mirror next to IM1 is for the green beams to be delivered to the PSL table. I think we still want to keep it. Otherwise, the plan looks fine. |
15654
|
Mon Nov 2 16:46:06 2020 |
gautam | Update | General | Plan for OMC chamber | To be a bit more clear about what we are going to do in the OMC chamber, I marked-up some photos, see Attachments #1 and #2.
- OM5 will be rotated to bring the IFO AS beam straight out without any splitting to the OMC.
- OMMT, OMC, DCPD, DCPD transimpedance amp, and all peripheral optics associated with these components, will be removed. Many of these components are mounted on a breadboard and so removing that breadboard will take care of it. These are marked with pink Xs.
I anticipate that after this work, the only components on the table will be
- IM1, to send the PSL beam to the IMC.
- OMs 5 and 6 to bring the IFO AS beam out onto the AP table (in principle, we could try and eliminate both these optics, if the AS beam happens to exit through one of the viewports cleanly, we will not have any intervening objects in the way once the OMC and peripherals are removed).
- MMT2 for mode-matching the IMC transmission to the interferometer mode.
Are we in agreement with this plan?
See #15656 for the updated photo |
Attachment 1: IMG_2318.JPG
|
|
Attachment 2: IMG_2332.JPG
|
|
15653
|
Mon Nov 2 12:44:25 2020 |
gautam | Update | General | ETMY and OMC chamber heavy doors off | [koji, rana, gautam]
This morning, we did the following;
- Removed the ETMY chamber heavy door. It is stored on the rack at the east end.
- Removed OMC chamber heavy door. It is smaller than the other chamber doors, so doesn't sit on the standard size rack we have. So it is laid flat, on a clean sheet of ameristat, on a cart next to the NE corner of the PSL table.
- After taking some photos of the chamber and making sure the position of the suspension tower was marked by some stops, we removed the ETMY cage and moved it to the cleanroom area. The optic was EQ stopped during the transport, and the OSEMs were removed. The ETMY Oplev HeNe was turned off and the PSL was shuttered to allow us to work without goggles.
- The broken off magnet was retrieved from inside the OSEM. The shadow sensor voltage recovered a value of ~800cts, which means the LED/PD pair on the UR OSEM seems to work fine, and it was in fact the magnet blocking the PD that was the issue.
- In the cleanroom area, we removed the optic from the wire loop and placed it on the magnet gluing fixture. The wire is intact (for now), so there is some hope of re-suspending it in the same loop.
The OSEMs remain in the EY vacuum chamber. The next set of steps are:
- Clean the EP30-2 residue from the broken magnet joint - this will require some scrubbing with an acetone soaked scrub or similar implement.
- Reglue the broken magnet.
We will most likely work on this tomorrow. At ~1615, I briefly opened the PSL shutter and tweaked the IMC alignment. We will almost certainly change the pointing into the IMC when we remove the old OMC and rebalance that table, so care should be taken when working on that...
Quote: |
- We are now ready to take the doors off.
|
|
Attachment 1: CB84FD62-C48B-414F-8825-50C9DE0B47CF.jpeg
|
|
15652
|
Thu Oct 29 19:30:56 2020 |
gautam | Update | General | Vent complete |
- Jam nuts - checked that none of the nuts move by hand, which according to Steve, is sufficient. I recorded photos of all of them here.
- Lab particle count - at the SP table, I measured 25,910 / cf @ 0.3 um and 1220 /cf @ 0.5 um. Steve's guidance is that the latter number should be less than 10,000, so no issues there.
- Valves - closed all the annuli off, and also closed VM1, VA6 and V1. The procedure calls for VM3 to be opened, presumably so that the RGA remains pumped, but I see no reason why we can't just leave that RGA volume valved off.
- Started letting N2 into the main volume at ~3pm local time, by opening VV1. Attachment #1 shows the valve configuration adopted for this vent. Went up to 25 torr and then switched over to the "Ultra Zero" instrument grade air cylinders.
- Aimed for 3-5 torr/min rate of pressure increase locally. The full vent trend can be seen in Attachment #2.
- Stopped letting air into the main volume when P1a pressure was 700 torr, at which point I disconnected the cylinders from the main volume, and left VV1 open so that the IFO equilibriates to the lab air pressure. I used 4 full cylinders of the instrument grade air, which is par for the course.
- Since we anticipate opening the ETMY and output optics chambers, I also vented those annuli. The final state of the valves as I am leaving it for the night is shown in Attachment #3.
- I re-aligned the IMC mirrors so that I could lock the IMC at low power once again. Indeed, IMC trans of ~1400 cts was realized (see Attachment #4), c.f. ~1500 earlier today, I thought this is fine and didn't optimize further. I think the policy is to not lock the IMC in air unless necessary, so I'm leaving the PSL shutter closed for the night.
- The ITMs, ETMs and BS were re-aligned such that the Oplev spots are centered. I can see some higher order resonances of the green beams. The AS spot is fringing (the two ND=0.5 filters on the AS CCD camera were removed for better visibility). So this is fine as far as alignment is concerned.
We are now ready to take the doors off. I've already done the basic prep work (loosened bolts, cleaned chamber, carts for tools, fresh ameristat on portable HEPAs etc).
Quote: |
If everything else looks good, I'll start letting the dry N2 into the main volume after lunch.
|
|
Attachment 1: ventValveStatus.png
|
|
Attachment 2: vent_Oct2020.jpg
|
|
Attachment 3: ventValveStatus_overnight.png
|
|
Attachment 4: lowPowerIMC_afterVent.png
|
|
15651
|
Thu Oct 29 12:43:35 2020 |
gautam | Update | General | Vent Prep |
- Oplev HeNe at ETMY was replaced, see here for my earlier discussion on this.
- I thought this is a good idea since we want the Oplev as a coarse reference and it'd not be ideal if this HeNe dies during the time the optic is out of the chamber.
- New HeNe had 2.8mW of power output as measured with Ophir power meter. This is in line with what is expected from these Lumentum heads.
- I labelled the head with the power output and today's date, re-aligned the Oplev reflection onto its QPD.
- After this work, the Y-arm could be locked without the huge angular fluctuations that was visible earlier, 👍 .
- GTRY anomaly
- I actually judged that there is no anomaly.
- The GTRY CDS indicator is actually quite useless - the ADC saturates at ~3500 cts (and not 32768 as you would expect from a 16 bit ADC but that's the well-known whitening filter saturation problem). This should be fixed, but this is a task for later.
- I measured with a DMM the voltage when the TEM00 is locked to the cavity and GTRY is 0.3 (the nominal value these days), the DC voltage was ~5.6 V. The prompt reflection from the ETM registers ~6.5 V DC. So the mode-matching isn't stellar, but this is again a known issue, and can be fixed later.
- Other pre-vent checks
- The Oplevs had drifted significantly, I re-centered ITMs, ETMs and BS after aligning the arm cavities and green beams in the POX/POY lock state. See Attachment #1.
- I locked the PRMI on carrier and used this config to re-center the PRM Oplev, see Attachment #2.
- No further action was taken regarding SRM oplev.
- I checked the ALS noise, see Attachment #3. The X arm ALS has excess noise at ~100 Hz that certainly wasn't there previosuly - sigh. There is nothing I can find about any changes made at EX in the elog.
- Updated the "DriftMon" values, though I guess we don't even really use this anymore these days?
- Re-relieved the IMC WFS offsets.
- Cut the input power to the IMC from 1.007W to 100.1 mW (both numbers measured with Ophir power meter).
- Replaced the 10% beamsplitter in the MCREFL path on the AS table with a Y1 HR mirror. Note that there is no beam on the IMC WFS in this configuration.
- Was able to lock the IMC on low power to a TEM00 mode -
need to set up the low power autolocker. The IMC autolocker is now set to the low power settings, and I've tested it locks a couple of times. Attachment #5 shows the low power lock in StripTool.
- Walked around and looked at all the bellows - the jam nuts are up against their stops, and I can't move them with my hands, so I think that's okay.
If everything else looks good, I'll start letting the dry N2 into the main volume after lunch.
Quote: |
Now the green transmissions are visible by the green PDs. Attachment 9 shows the trans and ref of each green beams with and without locking to TEM00. The questionable green TRY was ~0.3. If we compare this with the histrical data (Attachment 10), it is about 1/4 of the value in the past. It's not too crazy but still quite low.
|
BTW, nice video! @ Koji, How difficult was it to edit it into this form? |
Attachment 1: preVentAlignment.png
|
|
Attachment 2: PRMIcarr_PRMalignment.png
|
|
Attachment 3: ALS_ool.pdf
|
|
Attachment 4: lowPowerMC.png
|
|
15650
|
Thu Oct 29 09:50:12 2020 |
anchal | Summary | Calibration | Preliminary calibration measurement taken | I went to 40m yesterday at around 2:30 pm and Koji showed me how to acquire lock in different arms and for different lasers. Finally, we took a preliminary measurement of shaking the ETMX at some discrete frequencies and looking at the beatnote frequency spectrum of X-end laser's fiber-coupled IR and Main laser's IR pick-off.
Basic controls and measurement 101 at 40m
- I learned a few things from Koji about how to align the cavity mirrors for green laser or IR laser.
- I learned how to use ASS and how to align the green end laser to the cavity. I also found out about the window at ETMX chamber where we can directly see the cavity mode, cool stuff.
- Koji also showed me around on how to use diaggui and awggui for taking measurements with any of the channels.
Preliminary measurement for calibration scheme
We verified that we can send discrete frequency excitation signals to ETMX actuators directly and see a corresponding peak in the spectrum of beatnote frequency between fiber-coupled X-end IR laser and main laser IR pickoff.
- I sent excitation signal at 200 Hz, 250 Hz and 270 Hz at C1:SUS-ETMX_LSC_EXC channel using awggui with an amplitude of 100 cts and gain of 2.
- I measured corresponding peaks in the beatnote spectrum using diaggui.
- Page 1 shows the ASD data for the 4 measurements taken with Hanning window and averaging of 10.
- Page 2 shows close up Spectrum data for the 4 measurements taken with flattop window and averaging of 10.
- I converted this frequency signal into displacement by using conversion factor
or .
If full interferometer had been locked, we could have used the DARM error signal output to calibrate it against this measurement.
Data |
Attachment 1: PreliminaryCalibrationData.pdf
|
|
15649
|
Wed Oct 28 20:42:53 2020 |
Koji | Update | General | Vent Prep |
- Two arms / BS / PRM / SRM were aligned. (Attachment 1)
- IMC was aligned by WFS and the WFS offsets were offloaded.
- Suspension Status Snapshot (Attachment 2)
- Oplevs are aligned (Attachment 3)
- Xarm green was aligned in the daytime. Xarm green refl DC (C1:ALS-X_REFL_DC_OUTPUT) was 620 (aligned) ~1300 (drifted). When unlocked, it was 3750.
- Yarm green: I saw no flash. We don't have functional PZT alignment since the ASY M2 PZT got broken. I went to the Yend. Something funky is going on with the Yend green. I struggled to have any flash of the cavity. The apertures were not so precise. I finally got TEM00 locked, but the modematching seems exteremely low (like 1/1000?). Basically I saw no power reduction of the refl when the cavity is locked. So at least the cavity was locked but we might need to revisit when we open the chamber
==> Gautam thinks it was not like that. So he will check the green alignment tomorrow (Thu).
- Item checking: I familialized myself with the yend crane operation. Today I learned that there is a power switch on the wall (Attachment 4). The yend has two heavy door storages (Attachments 4/5). The slings to lift the heavy door are in the crane cabinet along with the y arm (Attachment 6). I didn't yet try to find the "hammer" to hit the door if the o-ring stuck too strong, although that's optional.
- We want to reduce the PSL power. But Gautam wants to use the arm locking with the nominal power, it will be done tomorrow by him.
- The last thing is to check the green trans power. I noticed that the green trans beams are blocked by an HWP for the BHD LO path on the PSL table. (Attachment 7)
The HWP was moved and the process was recorded in the movie (Attachment 8). The fiber output was monitored by the BHD DC (aka AS110 DC) with the AS path blocked. The fiber output of 22.6mV (offset -2.5mV) was improved tio 29.1mV after the HWP move and the alignment adjustment.
- Now the green transmissions are visible by the green PDs. Attachment 9 shows the trans and ref of each green beams with and without locking to TEM00. The questionable green TRY was ~0.3. If we compare this with the histrical data (Attachment 10), it is about 1/4 of the value in the past. It's not too crazy but still quite low.
At this point, I'm leaving the lab. All the suspensions (incl SRM) are aligned. PSL/GRX/GRY shutters were left open. |
Attachment 1: Screen_Shot_2020-10-28_at_19.42.27.png
|
|
Attachment 2: Screen_Shot_2020-10-28_at_19.42.38.png
|
|
Attachment 3: Screen_Shot_2020-10-28_at_20.42.26.png
|
|
Attachment 4: P_20201028_222839.jpg
|
|
Attachment 5: P_20201028_223020.jpg
|
|
Attachment 6: P_20201028_223212.jpg
|
|
Attachment 7: 2020-10-28-233038.jpeg
|
|
Attachment 8: HWP_moving.mp4
|
Attachment 9: Screenshot_from_2020-10-29_01-04-06.png
|
|
Attachment 10: Screenshot_from_2020-10-29_01-16-53.png
|
|
15648
|
Wed Oct 28 14:07:47 2020 |
gautam | Update | General | ISS checkout | Ok I was using the PD in the black mount because Rana recommended it a few weeks ago.
Regarding the M2ISS, I acquired the hardware from QIL some months ago, including a circuit board, and 2 PDs. These had LEMO outputs though (not BNC), and the mounts are not 4". These photodiodes are what I'm using as the airBHD DCPDs right now, and some photos are here - are these the photodiodes you mentioned? Or are there yet more M2ISS photodiodes? I remember Johannes had some custom mounts extruded to make them 4" high, do you mean those? Can I retrieve them his Cryo setup?
BTW, my elog scraping shows only one spectra from Stefan in the ATF elog, and the performance there is more like 1e-7/rtHz @ 100 Hz, and that’s using a dedicated high BW servo circuit, not the SR560. Am I just missing the measurement of 2e-8/rtHz?
Quote: |
that little PD in the black mount was never very good. The AD829 is not a good opamp for transimpedance and especially not good for low frequencies. Stefan Ballmer and I were able to get 2e-8 out of these (@100 Hz) many years ago.
I wonder if we have some of Zach's M2ISS photodetectors around, perhaps in QIL or Cryo. I doubt that any of them are in use now. Those had good performance nad BNC output.
|
|
15647
|
Wed Oct 28 14:01:03 2020 |
not gautam | Update | General | ISS checkout | that little PD in the black mount was never very good. The AD829 is not a good opamp for transimpedance and especially not good for low frequencies. Stefan Ballmer and I were able to get 2e-8 out of these (@100 Hz) many years ago.
I wonder if we have some of Zach's M2ISS photodetectors around, perhaps in QIL or Cryo. I doubt that any of them are in use now. Those had good performance nad BNC output. |
15646
|
Wed Oct 28 09:35:00 2020 |
Koji | Update | CDS | RFM errors | I'm starting the model restarts from remote. Then later I'll show up in the lab to do more hard resets.
==> It seems that the RFM errors are gone. Here are the steps.
- Shutdown all the watchdogs
- login to c1iscex. Shutdown all the realtime models: rtcds kill --all
- login to c1iscey. Shutdown all the realtime models: rtcds kill --all
- run scripts/cds/rebootC1LSC.sh on pianosa
- reboot c1iscex
- reboot c1isxey
- Wait until all the machines/models are up by the script
- restart c1iscex models
- restart c1iscey models
- some IPC errors are still visible on the CDS status screen. Lauch c1daf and c1oaf
|
Attachment 1: Screen_Shot_2020-10-28_at_10.06.00.png
|
|
15645
|
Tue Oct 27 23:47:53 2020 |
gautam | Update | General | ISS checkout | I wanted to look into the ISS situation. Some weeks ago, I found the PD that was previously used as the in-loop photodiode. I wanted to use this and measure the open-loop RIN at a few places (to see if there's any variation and also to check its functionality). However, I didn't get very far tonight - for a start, the PD height is 3" (while our beam height is 4" everywhere outside the vacuum), and I needed to put together a circuit to supply the 5V bias and +/- 15 V since the transimpedance is done on the head. I was only able to do a low-level functionality test tonight, checked that the DC voltage output varied linearly with the incident power (calibrated against an NF1611 photodiode, data will be put up later). I didn't get to measuring any noise performance - is an incandescent light bulb still shot noise limited at ~10 Hz < f < 10kHz? Some notes:
- The PD is DC coupled, and has a transimpedance of 1 kohm (inverting AD829 does the transimpedance).
- Probably a daughter board should be made that supplies the DC power voltages and rotues the output signal to something more convenient like a BNC connector. This daughter board can then also implement a DC coupled path (for monitoring) and AC coupled path (for servoing, fc to be determined).
- SR560 based ISS was implemented some years ago but I think the improvement was only seen above 100 Hz, and that too was marginal, the stabilized RIN was 10^-6 (monitored on an out-of-loop photodiode I think, but unsure). We'd probably want to aim for at least an order of magnitude better. Unclear at this point why more suppression wasn't possible back then, was it just insufficient loop gain, or was the sensing noise too high? To be investigated.
Unconnected to this work - this problem reared its ugly head again (i noticed it yesterday morning already actually). I don't have the energy to embark on a fix tonight, Koji is going to be in the lab all day tomorrow and so he will fix it. |
15644
|
Mon Oct 26 17:26:26 2020 |
gautam | Update | IOO | Excess laser freq noise investigation | Apart from the questionable wiring on the Acromags, one other important difference is in the way the connections were made between the old VME crates to the Eurocrate backplanes, and how we do it now. The thick cables had their sheilds connected to the eurocrate ground (or at least, there was a dedicated ground lug on those cables which we screwed on to the ground terminals on the Eurocrate backplanes). However, in our current configuration, we interface the Acromag ADCs and DACs to the backplane via these adaptor boards. The shields of the DSUB cables are presumably NOT connected to the Eurocrate grounds. This should also be investigated as one potential cause of the grounding issue - while on some of the Eurocrate modules, the P1/P2 connectors may have either the "A" or "C" row of connectors shorted to ground, some may not, and the TTFSS may suffer from such an issue?
Note that we have this problem in all of the slow machines that were upgraded to Acromag (if this turns out to be the issue).
Quote: |
In fact, the problem was the grounding issue (presumably on the IOO racks).
|
|
15643
|
Mon Oct 26 13:35:58 2020 |
Koji | Update | IOO | Excess laser freq noise investigation | In fact, the problem was the grounding issue (presumably on the IOO racks).
A temporary differential receiver at the TTFSS side was built using an SR560 and a few ponoma cables. This removed the structures ~850Hz.
The MC Servo Output was disconnected from the TTFSS box and monitored with SR785. The 850Hz structure was kept visible no matter what cables, including all the acromag DB cables, were removed. This made me suspicious about the measurement setup. The SR785 was connected to an AC power strip under the SP table and this was too far from the IOO rack.
The SR785 was connected to the AC power strip on 1X2, and now the difference becomes clear. No matter if the acromag cables are connected or not, the connection (particularly ground connection) between the MC servo module and the TTFSS box causes the MC servo output contaminated. (Comparison between Blue and Orange of Attachment #1). During the measurement, the EPICS switch for the fast path was disengaged (=no signal) and the VCO gain (...so called. It's just the MC Servo Gain) was set to be 0dB.
To test if the differential receiving of the MC Servo Output at the PSL helps to reduce this noise, I've built a simple (hacky) differential receiver using an SR560. (Attachment #2)
This kept the noise level same as the disconnected case (Comparison between Green and Orange of Attachment #1, I don't think the difference between them is not significant), while the IMC is locked as before.
Note that we can see that the 36kHz line was significantly reduced. Did we remove this annoying noise?
After talking with Gautam, we decided to leave this configuration while the SE-Diff cable was replaced with a more robust one. (See Attachment #3)
The PSL laser frequency performance was evakluated in the following two ways as we did last week:
1) Use the beat frequency of the free running PSL and the Y-end laser (Attachment #4). The PSL shutter was closed and thus the IMC was not locked.
2) Use the IMC MCF while the IMC was locked. (Attachment #5)
For both cases, the improvement was confirmed.
I also tried to check the reported issue by Gautam on this elog. He used 1Hz BW, but I cheated with 16Hz BW and 10x12.8kHz span PSDs. (Attachment #6)
For the measurement, IN1 GAIN of the IMC Servo was set to be 0dB and the OUT2 was switched to monitor the IN1 noise, while IN1 was terminated by a 50Ohm.
As I mentioned above, the AC power of SR785 was taken from a 1X2 power strip. Is this the reason for the power line forest look less severe compared to the previous case???
Anyway, I tried to use the same differential receiving technique (but with gain of x100) to see if this helps. The differential receiver helped to reduce the structure above 50kHz. The floor noise level was observed to be higher. I didn't pursue this any further, but the forest of the power line looked like a part of the measurement noise. This is indicative that the grounding condition on 1X2 is really not great and we need to review the configuration of the acromag grounding. |
Attachment 1: MC_Servo_Output.pdf
|
|
Attachment 2: 20201026135735_IMG_0175.jpg
|
|
Attachment 3: 20201026153435_IMG_0176.jpg
|
|
Attachment 4: Screen_Shot_2020-10-26_at_1.15.54_PM.png
|
|
Attachment 5: Screen_Shot_2020-10-26_at_1.35.19_PM.png
|
|
Attachment 6: MC_Servo_Error_Mon.pdf
|
|
15642
|
Fri Oct 23 19:01:57 2020 |
Koji | Summary | PEM | PSL Particle Counter kit removed from the table | The particle counter on the 40m PSL was removed. The package was made together with the OMC lab particle counter (see the packing list below).
The kit was picked up by Radhika for a python code to read out the numbers.
=== Packing List ===
- MET ONE 227A particle counter
- used at the 40m. It has the particle reading and the temperature reading.
- Power supply adapter (AC/DC) for 227A
- Caution: It is not compatible with GT-321.
- MET ONE GT-321
- I found another type of particle counter in West Bridge.
- Power supply adapter (AC/DC) for GT-321. (Labeled "for GT-321")
- Caution: It is not compatible with 227A.
- DB9 cable for GT-321
- Air Filter G3111
- When you run a particle counter attach this filter instead of the dust collecting cup to keep the air in take of the particle counter clean. This should keep the particle level down to zero.
|
Attachment 1: P_20201022_173529.jpg
|
|
Attachment 2: P_20201022_173419.jpg
|
|
15641
|
Fri Oct 23 16:41:06 2020 |
Koji | Update | IOO | Excess laser freq noise investigation | [Koji, Rana]
We wanted to track down the excess noise seen in MC_F and other places (see the previous report by Gautam)
Setup1: The IMC was locked and MC_F signal between 500 and 1500Hz was observed. The DTT template was saved as /users/Templates/MC/MCF_noise_201023.xml
- Suspected mech resonance/jitter coupled with clipping or any other imperfections. Poked the various optics and optomechanics on the table. Basically no change. If we tap the laser chassis and the optics close to the laser source, we occasionally unlocked the IMC
- When we touched (lifted) the Innolight controller box from the shelf, for the first time we saw a significant change in the shape of the noise spectrum. The peak around the 700Hz shited towards lower frequency by a few %. Other peaks have no obvious change in the shapes and the heights.
- While observing the MC_F signal on the laptop, we went to the back of the laser controller. Placing a hand close to the fan clearly changes the peak frequency lower. By temporarily disconnecting the fan from the power supply for a short moment, the 700Hz peak could be eliminated. We also tried to see the noise level with the slow thermal servo and diagnosis DB cable disconnected, but we didn't see any significant change of the noise level.
Setup 2: Using the ALS phase tracker, we can observe the relative freq noise of the PSL laser and the ETMY AUX laser without any servo involved. This way we can freely disconnect any cables from the lasers. The measurement template for DTT was saved as /users/Templates/ALS/Y_ALS_FINE_PHASE_OUT_102320.xml
- Noise spectrum before disconnecting the cable (REF0, RMS REF1)
- The Fast PZT input to the PSL was disconnected => This made all the peaks (including the 700Hz) disappeared (REF2, RMS REF3)
- The Fast PZT input was restored as before, then the chain was disconnected at the input of the HV PZT driver (Thorlabs) => Again, this made the peaks disappeared (REF4, RMS REF5)
- The chain was disconnected at the input of the TTFSS box => Again, this made the peaks disappeared (REF6, RMS REF7)
- Disconnected the demod input and the AO cables from the IMC servo board => This made the peaks came back (REF8)
- Disconnected all the input/peripheral cables from the IMC servo board except for the connection to the TTFSS box => Still the excess noise was observed (REF9)
- In addition to the above, the cable to the FSS box was disconnected but the ground was still touching the MC servo board => This made the peaks disappeared (REF10)
The conclusion is that the noise is injected from the main circuit of the IMC servo board.
Next time we will check if the backplane connection is doing something wrong. Also, we'll test if the presence of the RF signals does something bad to the IMC board via EMI and RFI.
We have reverted the connection and tested if we lock the IMC and Y arm. ==> We saw at least they were locked for a short period. The things are still stabilizing, but left them turned on so they keep trying to lock automatically for the night. |
Attachment 1: plot.pdf
|
|
|